(2asquared+3a-1)(a squared-4a)
Answer by lenny460(1073)

(2a^2 + 3a - 1)(a^2 - 4a)
Multiply using the FOIL Method:
2a^2 x a^2 = 2a^4
2a^2 x -4a = -8a^3
3a x a^2 = 3a^3
3a x -4a = -12a^2
-1 x a^2 = -1a^2
-1 x -4a = 4a
2a^4 - 8a^3 + 3a^3 - 12a^2 - 1a^2 + 4a
Combine like terms:
-8a^3 + 3a^3 = -5a^3
-12a^2 - 1a^2 = -13a^2
The Answer:
2a^4 - 5a^3 - 13a^2 + 4a